# Intersection cohomology without spaces

### v1

@@ Abstract
(3.1) $f_{vw}(t)=\sum_{i\geq 0}t^{i}\operatorname{dim}{\operatorname{H}}^{2i}(\overline{X_{w}})_{p}.$
@@
(3.2) $f_{FG}(t)=\sum_{i\geq 0}t^{i}\operatorname{dim}\operatorname{IH}^{2i}(\overline{\mathcal{O}_{G}})_{p}.$
In particular, taking $G=\{0\}$ and $F=\sigma$ so that $\overline{\mathcal{O}_{G}}=X(\sigma)$ and $\mathcal{O}_{F}=\{p\}$ is the unique fixed point, we have
@@
with the generating classes in degree 22, and the equivariant cohomology ${\operatorname{H}}_{T}(X)$ and equivariant intersection cohomology $\operatorname{IH}_{T}(X)$ are naturally $R$-modules. The equivariant cohomology of a torus orbit $\mathcal{O}$ is the quotient polynomial ring
@@
$\operatorname{IH}_{T}(X(\Sigma))\hookrightarrow\bigoplus_{\sigma\in\Sigma}\operatorname{IH}_{T}(X(\Sigma))_{\mathcal{O}_{\sigma}}.$
@@
A fundamental result of Bernstein and Lunts [11, Section 14] makes this possible. For simplicity, we assume that the new stratum $S_{x}=\{o\}$ is a single point fixed by $T$, and that there is an affine neighborhood NN of $o$ and a one-dimensional subtorus ${\mathbb{C}}^{*}\subset T$ which contracts NN onto $\{o\}$; see Remark 4.2 for a discussion of the more general case. Let $N^{\circ}:=N\setminus\{o\}$.
@@
where $T_{p}$ is the stabilizer of pp. We also assume that there exists a $T_{p}$-equivariant affine normal slice NN to $S_{x}$ at the point $p$, and that the stratum $S_{x}$ has a tubular neighborhood $\widehat{N}$ that is an NN-bundle over $S_{x}$. Letting $\widehat{N}^{\circ}=\widehat{N}\setminus S_{x}$, we have
@@
The sheaf $\mathcal{A}$ is flabby if and only if $\Sigma$ is a simplicial fan, meaning that each cone $\tau$ is generated by $\operatorname{dim}\tau$ rays. In that case, the equivariant intersection cohomology sheaf $\mathcal{F}$ coincides with the structure sheaf $\mathcal{A}$ itself. If $\Sigma$ is rational, this reflects the fact that the intersection cohomology of an orbifold coincides with its cohomology. For example, if $\Sigma$ is the fan in ${\mathbb{R}}^{2}$ whose maximal cones are the four orthants, and $R={\mathbb{R}}[x,y]$ is the polynomial ring associated with any of the maximal cones, then $\mathcal{F}(\Sigma)=\mathcal{A}(\Sigma)$ is the free $R$-module with basis consisting of the functions 11, $|x|$, $|y|$ and $|xy|$.
@@
For a non-simplicial example, we can lift the fan from the previous example to ${\mathbb{R}}^{3}$ by lifting the one-dimensional cones to the rays through $(1,0,1)$, (1,0,1)(-1,0,1), $(0,1,1)$, (0,1,1)(0,-1,1), and lifting the two-dimensional cones accordingly. The resulting fan is the boundary of a 33-dimensional cone $\hat{\sigma}$. Let $\widehat{\Sigma}$ be the fan consisting of $\hat{\sigma}$ and all of its faces, and let $\hat{\mathcal{F}}$ be the equivariant intersection cohomology sheaf on $\widehat{\Sigma}$.
@@
so $M_{\partial\hat{\sigma}}$ is the space of conewise polynomial functions on the boundary of $\hat{\sigma}$. By projecting down to ${\mathbb{R}}^{2}$, we see that this is isomorphic to the space of sections $\mathcal{F}(\Sigma)$ from the previous example, but now it is a module over the larger polynomial ring $\hat{R}={\mathbb{R}}[x,y,z]$. The multiplication by the new parameter $z$ is multiplication by $|x|+|y|$. It follows that $M_{\partial\hat{\sigma}}$ is a non-free $\hat{R}-$module with minimal generating set $\{1,|x|\}$, so the structure sheaf $\mathcal{A}$ of $\widehat{\Sigma}$ is not flabby. Instead, $M_{\sigma}$ is a free $\hat{R}-$module with generators in degree 0 and 22. This corresponds to the fact that the $g$-polynomial of the square is $g_{\square}(t)=1+t$.
@@
The last family of KLS-polynomials that we discuss is those coming from matroids, as in Example $3.$ Once again, we begin with the case of realizable matroids, in which there is a geometric object for us to study.
@@
Suppose that $|E|=1$ and $L={\mathbb{C}}$, so that $Y(L)={\mathbb{C}}\mathbb{P}^{1}$. The $R$-algebra ${\operatorname{H}}_{T}(Y(L))$ has an additive basis consisting of $y_{\emptyset}$ and $y_{E}$, where $y_{\emptyset}$ is the multiplicative identity and $y_{E}^{2}=\hbar y_{E}$. We begin with the isomorphisms $\mathcal{L}_{\emptyset}^{\circ}=\{E\}$ and $\mathcal{F}(\{E\})=\mathcal{A}(\{E\})=R$. We have two different homomorphisms from ${\operatorname{H}}_{T}(Y(L))$ to $R$: one is $\rho_{\emptyset}$, which takes $y_{\emptyset}$ to 11 and $y_{E}$ to 0; the other is the restriction map from $\mathcal{H}_{T}(\mathcal{L})$ to $\mathcal{H}_{T}(\{E\})$, which takes $y_{\emptyset}$ to 11 and $y_{E}$ to $\hbar$. We define
@@
More precisely, this is true provided that the graph is “$p$-GKM”, which means that the labels on edges adjacent to the same vertex do not become parallel modulo pp. For finite Weyl groups this rules out a few small primes, but for affine Weyl groups the restriction is more serious. The graph of a Schubert variety in the affine flag variety will be $p$-GKM for all but finitely many $p$, but the list of bad primes grows as the Schubert variety gets larger, and there are no characteristics that work for the graph of the entire affine flag variety.
@@
  • the lattice $\mathcal{L}$ is modular, and for every rank 22 interval $F,G]\subset\mathcal{L}$, we have $\{H\mid F<H<G\}\not\equiv 1\!\!\mod p$.

### v2

@@ Abstract
(3.1) $f_{vw}(t)=\sum_{i\geq 0}t^{i}\operatorname{dim}\operatorname{IH}^{2i}(\overline{X_{w}})_{p}.$
@@
(3.2) $f_{FG}(t)=\sum_{i\geq 0}t^{i}\operatorname{dim}\operatorname{IH}^{2i}(\overline{O_{G}})_{p}.$
In particular, taking $G=\{0\}$ and $F=\sigma$ so that $\overline{O_{G}}=X(\sigma)$ and $O_{F}=\{p\}$ is the unique fixed point, we have
@@
with the generating classes in degree $2$, and the equivariant cohomology ${\operatorname{H}}_{T}(X)$ and equivariant intersection cohomology $\operatorname{IH}_{T}(X)$ are naturally $R$-modules. The equivariant cohomology of a torus orbit $\mathcal{O}$ is the quotient polynomial ring
@@
$\operatorname{IH}_{T}(X(\Sigma))\hookrightarrow\bigoplus_{\sigma\in\Sigma}\operatorname{IH}_{T}(X(\Sigma))_{O_{\sigma}}.$
@@
A fundamental result of Bernstein and Lunts [11, Section 14] makes this possible. For simplicity, we assume that the new stratum $S_{x}=\{o\}$ is a single point fixed by $T$, and that there is an affine neighborhood $N$ of $o$ and a one-dimensional subtorus ${\mathbb{C}}^{*}\subset T$ which contracts $N$ onto $\{o\}$; see Remark 4.2 for a discussion of the more general case. Let $N^{\circ}:=N\setminus\{o\}$.
@@
where $T_{p}$ is the stabilizer of $p$. We also assume that there exists a $T_{p}$-equivariant affine normal slice $N$ to $S_{x}$ at the point $p$, and that the stratum $S_{x}$ has a tubular neighborhood $\widehat{N}$ that is an $N$-bundle over $S_{x}$. Letting $\widehat{N}^{\circ}=\widehat{N}\setminus S_{x}$, we have
@@
The sheaf $\mathcal{A}$ is flabby if and only if $\Sigma$ is a simplicial fan, meaning that each cone $\tau$ is generated by $\operatorname{dim}\tau$ rays. In that case, the equivariant intersection cohomology sheaf $\mathcal{F}$ coincides with the structure sheaf $\mathcal{A}$ itself. If $\Sigma$ is rational, this reflects the fact that the intersection cohomology of an orbifold coincides with its cohomology. For example, if $\Sigma$ is the fan in ${\mathbb{R}}^{2}$ whose maximal cones are the four orthants, and $R={\mathbb{R}}[x,y]$ is the polynomial ring associated with any of the maximal cones, then $\mathcal{F}(\Sigma)=\mathcal{A}(\Sigma)$ is the free $R$-module with basis consisting of the functions $1$, $|x|$, $|y|$ and $|xy|$.
@@
For a non-simplicial example, we can lift the fan from the previous example to ${\mathbb{R}}^{3}$ by lifting the one-dimensional cones to the rays through $(1,0,1)$, (1,0,1)(-1,0,1), $(0,1,1)$, (0,1,1)(0,-1,1), and lifting the two-dimensional cones accordingly. The resulting fan is the boundary of a $3$-dimensional cone $\hat{\sigma}$. Let $\widehat{\Sigma}$ be the fan consisting of $\hat{\sigma}$ and all of its faces, and let $\hat{\mathcal{F}}$ be the equivariant intersection cohomology sheaf on $\widehat{\Sigma}$.
@@
so $M_{\partial\hat{\sigma}}$ is the space of conewise polynomial functions on the boundary of $\hat{\sigma}$. By projecting down to ${\mathbb{R}}^{2}$, we see that this is isomorphic to the space of sections $\mathcal{F}(\Sigma)$ from the previous example, but now it is a module over the larger polynomial ring $\hat{R}={\mathbb{R}}[x,y,z]$. The multiplication by the new parameter $z$ is multiplication by $|x|+|y|$. It follows that $M_{\partial\hat{\sigma}}$ is a non-free $\hat{R}-$module with minimal generating set $\{1,|x|\}$, so the structure sheaf $\mathcal{A}$ of $\widehat{\Sigma}$ is not flabby. Instead, $M_{\sigma}$ is a free $\hat{R}-$module with generators in degree 0 and $2$. This corresponds to the fact that the $g$-polynomial of the square is $g_{\square}(t)=1+t$.
@@
The last family of KLS-polynomials that we discuss is those coming from matroids, as in Example 3.3. Once again, we begin with the case of realizable matroids, in which there is a geometric object for us to study.
@@
Suppose that $|E|=1$ and $L={\mathbb{C}}$, so that $Y(L)={\mathbb{C}}\mathbb{P}^{1}$. The $R$-algebra ${\operatorname{H}}_{T}(Y(L))$ has an additive basis consisting of $y_{\emptyset}$ and $y_{E}$, where $y_{\emptyset}$ is the multiplicative identity and $y_{E}^{2}=\hbar y_{E}$. We begin with the isomorphisms $\mathcal{L}_{\emptyset}^{\circ}=\{E\}$ and $\mathcal{F}(\{E\})=\mathcal{A}(\{E\})=R$. We have two different homomorphisms from ${\operatorname{H}}_{T}(Y(L))$ to $R$: one is $\rho_{\emptyset}$, which takes $y_{\emptyset}$ to $1$ and $y_{E}$ to 0; the other is the restriction map from $\mathcal{H}_{T}(\mathcal{L})$ to $\mathcal{H}_{T}(\{E\})$, which takes $y_{\emptyset}$ to $1$ and $y_{E}$ to $\hbar$. We define
@@
More precisely, this is true provided that the graph is “$p$-GKM”, which means that the labels on edges adjacent to the same vertex do not become parallel modulo $p$. For finite Weyl groups this rules out a few small primes, but for affine Weyl groups the restriction is more serious. The graph of a Schubert variety in the affine flag variety will be $p$-GKM for all but finitely many $p$, but the list of bad primes grows as the Schubert variety gets larger, and there are no characteristics that work for the graph of the entire affine flag variety.
@@
  • the lattice $\mathcal{L}$ is modular, and for every rank $2$ interval $F,G]\subset\mathcal{L}$, we have $\{H\mid F<H<G\}\not\equiv 1\!\!\mod p$.